# Anabelian aspects of the outer automorphism groups of the absolute Galois groups of mixed-characteristic local fields

### v1

@@ 5. $(d)$
    The centralizer of $H\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\operatorname{Aut}(k)\subset\operatorname{Aut}_{\mathbb{Q}_{p_{k}}}(k_{+})=\operatorname{GL}_{d_{k}}(\mathbb{Q}_{p_{k}})$ in $\operatorname{GL}_{d_{k}}(\mathbb{Q}_{p_{k}})$ is a closed subgroup of $\operatorname{GL}_{d_{k}}(\mathbb{Q}_{p_{k}})$. In particular, we regard the centralizer of $H$ in $\operatorname{GL}_{d_{k}}(\mathbb{Q}_{p_{k}})$ as a $p_{k}$-adic Lie subgroup of $\operatorname{GL}_{d_{k}}(\mathbb{Q}_{p_{k}})$ [cf. [D-S-M-S], Theorem 9.6].
@@
which is induced by standard generating loops of $S\setminus{P}$ [cf., e.g., [Hat], the discussion following Proposition 1.26 and related arguments]. Define $\delta\stackrel{{\scriptstyle\mathrm{def}}}{{=}}[a_{1},b_{1}]\cdots[a_{g},b_{g}]=c^{-1}$ and $\mathrm{Aut}^{\delta}(\pi_{1}(S\setminus\{P\}))\subset\operatorname{Aut}(\pi_{1}(S\setminus\{P\}))$ to be the subgroup of automorphisms of $\pi_{1}(S\setminus\{P\})$ that fix $\delta\in\pi_{1}(S\setminus\{P\})$. For each integer $i$ satisfying $1\leq i\leq g$, we assign
@@
Let $\operatorname{Mod}(S\setminus\{P\})$ be the mapping class group of $S\setminus\{P\}$ [see [F-M], the beginning of Chapter 2]. Consider the natural homomorphism $\operatorname{Mod}(S\setminus\{P\})\to\operatorname{Out}(\pi_{1}(S\setminus\{P\}))$. It is well-known that the natural map $\operatorname{Mod}(S\setminus\{P\})\to\operatorname{Out}(\pi_{1}(S\setminus\{P\}))$ is injective, and that there exists a lifting in $\operatorname{Aut}(\pi_{1}(S\setminus\{P\}))$ of the image of each element of $\operatorname{Mod}(S\setminus\{P\})$ which fixes $\delta$ [cf. [F-M], Theorem 4.1, and the remark following [F-M], Theorem 4.1]. Throughout the remainder of the present paper, we regard $\operatorname{Mod}(S\setminus\{P\})$ as a subgroup of $\operatorname{Out}(\pi_{1}(S\setminus\{P\}))$ via the natural injective group homomorphism $\operatorname{Mod}(S\setminus\{P\})\hookrightarrow\operatorname{Out}(\pi_{1}(S\setminus\{P\}))$. Let $\varphi$ be an element of $\operatorname{Mod}(S\setminus\{P\})$. As mentioned above, there exists a lifting of $\varphi$ to $\operatorname{Aut}(\pi_{1}(S\setminus\{P\}))$ which fixes $\delta$. Hence, each such lifting induces an automorphism of $G_{k}$.
@@
Let us consider the automorphisms $\varphi_{i}$, $\varphi^{\prime}_{i}$, and $\varphi^{\prime\prime}_{i}$ of $G_{k}$, which are defined in the proof of Theorem 1.3. Then we observe that, in light of the relation of $\operatorname{Out}(G_{k})$ with $\operatorname{Mod}(S\setminus\{P\})$ discussed above, for each integer $1\leq i\leq g$, $\varphi_{i}$ corresponds to the Dehn twist determined by a longitude, $\varphi^{\prime}_{i}$ corresponds to the Dehn twist determined by a meridian, and, if $g\geq 2$, then the element $\varphi^{\prime\prime}_{i}$ corresponds to the Dehn twist determined by a loop spanning two adjacent holes of a torus [cf. [Kum], Lemma 7.3.4].
@@
This assertion follows from the naturality of the isomorphism $\pi_{1}(S\setminus\{P\})^{\operatorname{ab}}\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}H_{1}(S\setminus\{P\},\mathbb{Z})$; [F-M], Theorem 6.4; and the remark following [F-M], Theorem 6.4. ∎
@@
We first consider the case where $g\geq 2$. It follows from Lemma 2.13 that the map $\Phi$ [cf. Remark 2.18] is a closed map. Thus, both $I\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\Phi(\operatorname{Out}(G_{k}))$ and $Z\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\Phi(Z_{\operatorname{Out}(G_{k})}(\operatorname{Aut}(k)))$ are closed subgroups of $\operatorname{GL}_{d_{k}}(\mathbb{Q}_{p_{k}})$. In particular, it follows from [D-S-M-S], Theorem 9.6, that both $I$ and $Z$ naturally admit structures of $p_{k}$-adic Lie groups. Furthermore, it follows from Remark 2.18 that the image of $D$ [cf. the discussion preceding Remark 2.16] by $\Phi$ is equal to $\mathrm{Sp}_{2g}(\mathbb{Z}_{p_{k}})$, which implies that $\mathrm{Sp}_{2g}(\mathbb{Z}_{p_{k}})\subset I$. As is well-known, since the dimension of $\mathrm{Sp}_{2g}(\mathbb{Z}_{p_{k}})$ as a $p_{k}$-adic Lie group is $2g^{2}+g$, it follows that $\dim(I)\geq 2g^{2}+g$.
@@
When $p_{k}=2$ and $\sqrt{-1}\in k$, as in the case where $p_{k}$ is odd, an explicit description by generators and relations for the topological group $G_{k}$ [as discussed in Theorem 1.1] is known. In fact, in [Nis], Nishio proved similar results to Theorem 1.9 and Theorem 2.9 in the situation where $k$ is absolutely abelian, and $\sqrt{-1}\in k$ by means of the explicit description. On the other hand, when $p_{k}=2$ and $\sqrt{-1}\notin k$, a similar explicit description by generators and relations for the topological group $G_{k}$ is not known. In particular, it is not clear to the author at the time of writing whether or not the subgroup $\operatorname{Aut}(k)\subset\operatorname{Out}(G_{k})$ is a normal subgroup in the case where $p_{k}=2$ and $\sqrt{-1}\notin k$, even when $k$ is absolutely Galois.

### v2

@@ 5. $(d)$
    The centralizer of $H\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\operatorname{Aut}(k)\subset\operatorname{Aut}_{\mathbb{Q}_{p_{k}}}(k_{+})=\operatorname{GL}_{d_{k}}(\mathbb{Q}_{p_{k}})$ in $\operatorname{GL}_{d_{k}}(\mathbb{Q}_{p_{k}})$ is a closed subgroup of $\operatorname{GL}_{d_{k}}(\mathbb{Q}_{p_{k}})$. In particular, we regard the centralizer of $H$ in $\operatorname{GL}_{d_{k}}(\mathbb{Q}_{p_{k}})$ as a $p_{k}$-adic Lie subgroup of $\operatorname{GL}_{d_{k}}(\mathbb{Q}_{p_{k}})$ [cf. [1], Theorem 9.6].
@@
which is induced by standard generating loops of $S\setminus{P}$ [cf., e.g., [5], the discussion following Proposition 1.26 and related arguments]. Define $\delta\stackrel{{\scriptstyle\mathrm{def}}}{{=}}[a_{1},b_{1}]\cdots[a_{g},b_{g}]=c^{-1}$ and $\mathrm{Aut}^{\delta}(\pi_{1}(S\setminus\{P\}))\subset\operatorname{Aut}(\pi_{1}(S\setminus\{P\}))$ to be the subgroup of automorphisms of $\pi_{1}(S\setminus\{P\})$ that fix $\delta\in\pi_{1}(S\setminus\{P\})$. For each integer $i$ satisfying $1\leq i\leq g$, we assign
@@
Let $\operatorname{Mod}(S\setminus\{P\})$ be the mapping class group of $S\setminus\{P\}$ [see [2], the beginning of Chapter 2]. Consider the natural homomorphism $\operatorname{Mod}(S\setminus\{P\})\to\operatorname{Out}(\pi_{1}(S\setminus\{P\}))$. It is well-known that the natural map $\operatorname{Mod}(S\setminus\{P\})\to\operatorname{Out}(\pi_{1}(S\setminus\{P\}))$ is injective, and that there exists a lifting in $\operatorname{Aut}(\pi_{1}(S\setminus\{P\}))$ of the image of each element of $\operatorname{Mod}(S\setminus\{P\})$ which fixes $\delta$ [cf. [2], Theorem 4.1, and the remark following [2], Theorem 4.1]. Throughout the remainder of the present paper, we regard $\operatorname{Mod}(S\setminus\{P\})$ as a subgroup of $\operatorname{Out}(\pi_{1}(S\setminus\{P\}))$ via the natural injective group homomorphism $\operatorname{Mod}(S\setminus\{P\})\hookrightarrow\operatorname{Out}(\pi_{1}(S\setminus\{P\}))$. Let $\varphi$ be an element of $\operatorname{Mod}(S\setminus\{P\})$. As mentioned above, there exists a lifting of $\varphi$ to $\operatorname{Aut}(\pi_{1}(S\setminus\{P\}))$ which fixes $\delta$. Hence, each such lifting induces an automorphism of $G_{k}$.
@@
Let us consider the automorphisms $\varphi_{i}$, $\varphi^{\prime}_{i}$, and $\varphi^{\prime\prime}_{i}$ of $G_{k}$, which are defined in the proof of Theorem 1.3. Then we observe that, in light of the relation of $\operatorname{Out}(G_{k})$ with $\operatorname{Mod}(S\setminus\{P\})$ discussed above, for each integer $1\leq i\leq g$, $\varphi_{i}$ corresponds to the Dehn twist determined by a longitude, $\varphi^{\prime}_{i}$ corresponds to the Dehn twist determined by a meridian, and, if $g\geq 2$, then the element $\varphi^{\prime\prime}_{i}$ corresponds to the Dehn twist determined by a loop spanning two adjacent holes of a torus [cf. [12], Lemma 7.3.4].
@@
This assertion follows from the naturality of the isomorphism $\pi_{1}(S\setminus\{P\})^{\operatorname{ab}}\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}H_{1}(S\setminus\{P\},\mathbb{Z})$; [2], Theorem 6.4; and the remark following [2], Theorem 6.4. ∎
@@
We first consider the case where $g\geq 2$. It follows from Lemma 2.13 that the map $\Phi$ [cf. Remark 2.18] is a closed map. Thus, both $I\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\Phi(\operatorname{Out}(G_{k}))$ and $Z\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\Phi(Z_{\operatorname{Out}(G_{k})}(\operatorname{Aut}(k)))$ are closed subgroups of $\operatorname{GL}_{d_{k}}(\mathbb{Q}_{p_{k}})$. In particular, it follows from [1], Theorem 9.6, that both $I$ and $Z$ naturally admit structures of $p_{k}$-adic Lie groups. Furthermore, it follows from Remark 2.18 that the image of $D$ [cf. the discussion preceding Remark 2.16] by $\Phi$ is equal to $\mathrm{Sp}_{2g}(\mathbb{Z}_{p_{k}})$, which implies that $\mathrm{Sp}_{2g}(\mathbb{Z}_{p_{k}})\subset I$. As is well-known, since the dimension of $\mathrm{Sp}_{2g}(\mathbb{Z}_{p_{k}})$ as a $p_{k}$-adic Lie group is $2g^{2}+g$, it follows that $\dim(I)\geq 2g^{2}+g$.
@@
When $p_{k}=2$ and $\sqrt{-1}\in k$, as in the case where $p_{k}$ is odd, an explicit description by generators and relations for the topological group $G_{k}$ [as discussed in Theorem 1.1] is known. In fact, in [17], Nishio proved similar results to Theorem 1.9 and Theorem 2.9 in the situation where $k$ is absolutely abelian, and $\sqrt{-1}\in k$ by means of the explicit description. On the other hand, when $p_{k}=2$ and $\sqrt{-1}\notin k$, a similar explicit description by generators and relations for the topological group $G_{k}$ is not known. In particular, it is not clear to the author at the time of writing whether or not the subgroup $\operatorname{Aut}(k)\subset\operatorname{Out}(G_{k})$ is a normal subgroup in the case where $p_{k}=2$ and $\sqrt{-1}\notin k$, even when $k$ is absolutely Galois.